# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 3.2 Physics-based Methods
Physics-based methods accomplish motion prediction by simulating a set of dynamics equations describing a physics-based model. These methods apply the laws of physics to estimate steering and acceleration, predicting position, orientation, and speed. To describe vehicle movement, they employ kinematic [204] and dynamic models. Kinematic models, which do not account for the forces governing motion, are widely used due to their simplicity. Commonly used kinematic models include the Constant-Velocity (CV) model, which estimates position and velocity under the assumption of constant velocity; the Constant-Acceleration (CA) model, which estimates position, velocity, and acceleration assuming constant acceleration; and the Constant Turn Rate and Velocity (CTRV) and Constant Turn Rate and Acceleration (CTRA) models, which estimate position and orientation under the assumption of constant yaw rate. Dynamic models, on the other hand, account for forces impacting motion, introducing significant complexity when modeling components such as wheels, engines, and gearboxes. However, the added complexity in dynamic models often results in only marginal improvements in prediction accuracy. Furthermore, the forces governing the motion of other vehicles are not observable through sensors. These limitations make dynamic models less preferable than kinematic models, with a tendency toward simpler dynamic models for practical applications.
@@
The simplest method to predict a trajectory is to directly apply the vehicle’s current state to dynamic [205, 206, 207, 208] or kinematic models [209, 210]. This method assumes no noise in the vehicle’s state estimation, making it suitable for short-term prediction. However, as it does not account for state uncertainty, it becomes unreliable for long-term motion prediction. Kalman Filtering (KF) Methods address state uncertainty by representing the vehicle state with a Gaussian distribution. The state includes position $(x,y)$ and velocity $(v_{x},v_{y})$, and its prediction is computed from the current state using a state transition matrix defined by dynamic or kinematic models. The filter updates the covariance matrix of the state estimation error based on system dynamics, predicts the expected measurement, and computes the error between predicted and actual measurements. The Kalman gain is then calculated to update the state estimate and reduce uncertainty. Variants of Kalman filters, such as the linear Kalman Filter [211], the Extended Kalman Filter (EKF) [212], and the Unscented Kalman Filter (UKF) [213], have been applied to non-linear systems in trajectory prediction.
@@
While KF methods assume Gaussian distributions, Monte Carlo Methods approximate state distributions without such assumptions. These methods randomly sample input variables (e.g., initial conditions, sensor noise) and apply a physics model to generate potential future trajectories [214]. Monte Carlo methods are effective in estimating uncertainty but are computationally intensive and require a well-defined set of parameters and constraints to generate feasible results. Several works extend physics-based methods by embedding map information and incorporating road constraints. Petrich et al. [215] proposed a stochastic filter that uses digital map information to select a set of reasonable trajectories by projecting detected vehicles onto the centerlines of map lanes. Rehder and Kloeden [216] mapped environments to location maps representing pedestrian locations from occupancy grids of sidewalks and obstacles, converting the prediction problem into a planning task by modelling pedestrian destinations as latent variables. Coscia et al. [93] used polar grids centered at predicted agents to construct point-wise circular distributions, combining them into statistically smoothed predictions. Next, Koschi et al. [217] incorporated pedestrian dynamics, traffic rules, road connectivity, and semantic regions to obtain reachable pedestrian occupancy. Batkovic et al. [218] employed graph representations of road geometry to predict pedestrian motion, assuming motion along graph edges representing sidewalks and crosswalks, with a unicycle model ensuring trajectory adherence to edges.
@@
In real traffic, agent motion behavior is complex and cannot be described by a single model, often involving linear movements, turns, and acceleration/deceleration modes. Since motion modes of other vehicles are not directly observable, Multiple-Model (MM) methods address this uncertainty by using a set of possible motion models. These methods employ a bank of elemental filters, each based on a model from the set, to represent complex motion. At each time step, candidate models are evaluated, and the filters generate estimates, which can be combined or fused. For further details on handling discrete and continuous-based uncertainties, the review on MM methods [219] provides comprehensive insights. The Interacting Multiple Model (IMM) estimator [220, 221] is a hybrid filter with nearly linear computational complexity, offering a balance between complexity and performance. Kaempchen et al. [222] proposed an IMM approach utilizing constant speed, constant acceleration, and acceleration change models to compute vehicle states as weighted state estimates. Other IMM designs accounting for road constraints are presented in [223, 224]. Recent works, such as Interacting Multiple Model Kalman Filter (IMM-KF) [225], employs intention-based models and optimization-based projections for generating non-colliding predictions using a hierarchical approach for multiple vehicles. Kuhnt et al. [226] proposed a Unified Traffic Situation Estimation Model using IMM with constant and target velocity models, incorporating predefined environmental geometry to estimate future trajectories.
@@
Physics-based models are efficient for short-term predictions, but for long-term predictions, hybrid models [227, 228, 229, 230] have been developed by combining physics-based and learning-based approaches. Kim et al. [227] proposed a hybrid method that uses the Reciprocal Velocity Obstacles (RVO) framework to predict pedestrian trajectories in dynamic environments, enhanced with Ensemble Kalman Filtering (EnKF) and Expectation-Maximization (EM) for real-time learning and parameter refinement. Ma et al. [228] introduced a deep latent variable model built on a Wasserstein auto-encoder, enabling the automatic satisfaction of kinematic constraints. Yue et al. [229] presented Neural Social Physics (NSP), a Neural Differential Equation model that integrates an explicit physics-based framework as an inductive bias, paired with deep neural networks for system parameter estimation and stochastic dynamics modelling.
@@ 3.3 Machine Learning-based Methods
Machine Learning-based methods learn motion dynamics from driving data by fitting function approximators. Figure 9 provides an overview of two key ML techniques discussed in this section: Gaussian Processes (GPs) and Hidden Markov Models (HMMs). Additionally, the section discusses prediction models based on Gaussian Mixture Models (GMMs), which combine multiple GPs, and Dynamic Bayesian Networks (DBNs), which effectively capture interactions between agents. Each method is discussed in dedicated subsections, elaborating on the theoretical foundations, their application to prediction tasks, and relevant studies available in the literature.
@@ 3.3.2 Hidden Markov Models
HMM-based models are particularly effective in scenarios where the agent’s behavior is governed by discrete modes or intentions. Rathore et al. [241] proposed a hybrid framework, Traj-clusiVAT, combining trajectory clustering and Markov chain modelling. The clustering algorithm partitions trajectories into representative clusters, and Markov models are trained for prediction. Li et al. [242] proposed an enhanced HMM for behavior prediction, with key improvements including random initial parameter generation to ensure reliability, parameter learning through the Baum-Welch algorithm, and prediction via the Viterbi algorithm combined with Bayesian decision theory. On Continuous HMMs, Ren et al. [243] proposed a framework for behavior recognition. The CHMM is employed for early detection of lane-changing behavior by analyzing vehicle motion patterns and contextual traffic features.
@@ 3.4 Deep Learning-based Methods
Deep Learning-based methods excel in learning motion dynamics from driving data and are particularly effective for long-term predictions in complex environments. Figure 10 illustrates the main DL models reviewed in this section. Each approach is detailed in a separate subsection, with methods summarized in tables that highlight the key features of the architectures.
3.4.1 Sequential Models
@@
Sequential models operate under the assumption that motion can be described by temporally causal dependencies. To capture complex long-term temporal relationships, recurrent neural networks (RNNs) are widely used for trajectory prediction. RNNs and their variants effectively capture the dynamic properties of motion and the temporal correlations in prediction sequences by incorporating information from previous timestamps and using hidden states to generate future trajectory sequences. Focusing on model design, earlier studies primarily utilized a single recurrent model to map past trajectories to future paths [288, 97, 102]. However, more recent works employ multiple groups of recurrent models to address different aspects of motion [289, 256, 22], as well as stacked LSTMs and convolutional LSTMs. Dai et al. [22] proposed a model, where one group of LSTMs models individual vehicle motions, while another group captures interactions between agents. The architectural approach for stacked LSTMs involves hierarchically organizing recurrent networks into layers, as seen in [290, 251]. Tables 3 and 4 provide a summary of architectures in the literature that rely on recurrent models.
@@ 3.4.2 CNNs
Convolution prediction models are categorized into two groups in the literature: those using image-based input and Temporal Convolution Networks (TCNs). Image-based inputs can come from various perspectives, including BEV, vehicle-mounted cameras, or other sensor modalities. BEV images or aerial photographs offer a top-down perspective that effectively captures spatial relationships and road structures. CNN-based models [69, 61, 298] primarily focus on the feature extraction stage to capture road geometries, traffic agent positions, and environmental context. The second category involves Temporal Convolution Networks (TCNs), which apply convolution layers directly to trajectory sequences to model temporal dependencies. Unlike RNNs, TCNs operate in parallel over the entire sequence, incorporating causal convolutions to preserve temporal order and dilated convolutions to extend the receptive field, capturing long-term dependencies. TCNs are designed to maintain stable gradients, avoid information leakage, and achieve computational efficiency for long sequences. One example is the framework proposed by Nikhil and Morris [86], which preserves spatio-temporal consistency through multiple convolution layers while leveraging parallelism for efficiency. Table 5 provides a summary of CNN-based architectures for trajectory prediction.
@@ 3.4.3 Autoencoders
In trajectory prediction, autoencoders encode past trajectories into compact latent representations, which are then decoded to predict future positions. These models excel at capturing complex temporal dependencies and can function as standalone predictors or feature extractors integrated into larger architectures. Variational Autoencoders (VAEs) extend this approach by modeling uncertainty, generating a latent distribution instead of fixed representations. By sampling from this distribution, VAEs generate diverse and multimodal future trajectories, modelling the inherent uncertainty in trajectory forecasting tasks. Table 6 provides a summary of autoencoder architectures available in the literature, grouped by deterministic or stochastic type.
@@ 3.4.4 GANs
Generative Adversarial Networks (GANs) have emerged as a powerful framework for trajectory prediction, leveraging their ability to model complex, multimodal data distributions. In this context, the generator ($G$) predicts plausible future trajectories conditioned on past trajectories, while the discriminator ($D$) evaluates these predictions against real trajectory data to guide the generator’s learning. This adversarial training setup enables GANs to generate diverse and realistic trajectories that account for the inherent uncertainty in future motion. A key advantage of GAN-based models is their capacity to address the multimodal nature of trajectory prediction, where multiple plausible futures exist for a given past trajectory. This is achieved by conditioning $G$ on contextual information such as social interactions, environmental constraints, or road semantics, allowing it to adapt predictions to dynamic and complex scenarios. Table 7 provides a summary of GAN architectures in the literature, detailing the design of $G$ and $D$.
@@ 3.4.6 Attention and Transformers
Attention mechanisms enhance prediction models by allowing them to focus on the most relevant spatial and temporal features in complex environments. Spatial attention mechanisms assign importance weights to neighboring agents computed at each timestamp, based on features such as relative distance, velocity, and heading, which influence the target agent’s trajectory. Temporal attention processes sequences of historical states, assigning weights to time steps that have the most significant impact on future predictions. This is often integrated with recurrent architectures, where the hidden states from prior time steps are passed through an attention module to refine the context vector for trajectory prediction. Advanced implementations combine spatial and temporal attention in multi-head architectures, enabling the model to process multiple attention patterns in parallel, effectively handling heterogeneous agents in dense environments. Table 11 provides a summary of reviewed models relying on attention mechanisms.
@@ 3.5 Reinforcement Learning-based Methods
Reinforcement Learning-based methods generate predictions by interpreting agent behaviour as a high-dimensional policy optimization problem. RL methods learn through a reward system to generate trajectories that maximize long-term objectives. This section reviews two main categories of methods existing in the literature: Inverse Reinforcement Learning (IRL) and Imitation Learning.

### v2

@@ 3.2 Physics-based Methods
Physics-based methods accomplish motion prediction by simulating a set of dynamics equations describing a physics-based model. These methods apply physical laws to estimate steering and acceleration, predicting position, orientation, and speed. To describe vehicle movement, they employ kinematic [201] and dynamic models. Commonly used kinematic models include the Constant-Velocity (CV) model, which estimates position and velocity under the assumption of constant velocity; the Constant-Acceleration (CA) model, which estimates position, velocity, and acceleration assuming constant acceleration; and the Constant Turn Rate and Velocity (CTRV) and Constant Turn Rate and Acceleration (CTRA) models, which estimate position and orientation under the assumption of constant yaw rate. Dynamic models account for forces acting on the vehicle, which introduces added complexity when modeling components such as wheels, engines, and gearboxes. These limitations make dynamic models less preferable than kinematic models, with a tendency toward simpler dynamic formulations in practical applications.
@@
The simplest approach to trajectory prediction is to apply the vehicle’s current state to a dynamic model [202, 203, 204, 205] or a kinematic model [206, 207]. This approach assumes perfect state estimation and does not account for uncertainty. Kalman Filtering (KF) Methods address this limitation by modeling the vehicle state as a Gaussian distribution. The state typically includes position $(x,y)$ and velocity $(v_{x},v_{y})$, and its future estimate is computed using a state transition matrix defined by the underlying motion model. The filter updates the covariance of the state estimation error based on system dynamics, predicts the expected measurement, and corrects the state estimate using the Kalman gain derived from the prediction-error residual. Variants of Kalman Filters, such as the standard linear Kalman Filter [208], the Extended Kalman Filter (EKF) [209], and the Unscented Kalman Filter (UKF) [210] — have been employed for trajectory prediction in both linear and nonlinear systems.
@@
While KF methods assume Gaussian distributions, Monte Carlo Methods approximate state distributions without such assumptions. These methods randomly sample input variables (e.g., initial conditions, sensor noise) and apply a physics model to generate potential future trajectories [211]. Several works extend physics-based methods by embedding map information and incorporating road constraints. Petrich et al. [212] proposed a stochastic filter that uses digital map information to select a set of reasonable trajectories by projecting detected vehicles onto the centerlines of map lanes. Rehder and Kloeden [213] mapped environments to location maps representing pedestrian locations from occupancy grids of sidewalks and obstacles, converting the prediction problem into a planning task by modelling pedestrian destinations as latent variables. Coscia et al. [93] used polar grids centered at predicted agents to construct point-wise circular distributions, combining them into statistically smoothed predictions. Next, Koschi et al. [214] incorporated pedestrian dynamics, traffic rules, road connectivity, and semantic regions to obtain reachable pedestrian occupancy. Batkovic et al. [215] employed graph representations of road geometry to predict pedestrian motion, assuming motion along graph edges representing sidewalks and crosswalks, with a unicycle model ensuring trajectory adherence to edges.
@@
In real traffic, agent motion behavior is complex and cannot be described by a single model, often involving linear movements, turns, and acceleration/deceleration modes. Since motion modes of other vehicles are not directly observable, Multiple-Model (MM) methods address motion uncertainty by considering a set of predefined motion models. A bank of elemental filters is used, each corresponding to one model in the set. At each time step, the filters independently generate state estimates, and their outputs are evaluated and combined typically through weighted fusion to represent complex or switching motion patterns. For further details on handling discrete and continuous-based uncertainties, the review on MM methods [216] provides comprehensive insights. The Interacting Multiple Model (IMM) estimator [217, 218] is a hybrid filter with nearly linear computational complexity, offering a balance between complexity and performance. Kaempchen et al. [219] proposed an IMM approach utilizing constant speed, constant acceleration, and acceleration change models to compute vehicle states as weighted state estimates. Other IMM designs accounting for road constraints are presented in [220, 221]. Recent works, such as Interacting Multiple Model Kalman Filter (IMM-KF) [222], employs intention-based models and optimization-based projections for generating non-colliding predictions using a hierarchical approach for multiple vehicles. Kuhnt et al. [223] proposed a Unified Traffic Situation Estimation Model using IMM with constant and target velocity models, incorporating predefined environmental geometry to estimate future trajectories.
@@
3.2.1 Strengths and Limitations of Physics-based Models
Physics-based models are efficient for short-term trajectory prediction, particularly in situations requiring rapid adaptation to abrupt or unexpected maneuvers. Their dependence on current or recent vehicle state makes them computationally lightweight, but also limits their ability to incorporate longer motion histories. Constant motion models do not account for uncertainty, whereas Kalman Filter variants introduce Gaussian noise modeling but remain limited in flexibility. Monte Carlo methods offer greater expressiveness for modeling uncertainty without relying on specific distributional assumptions, but they are computationally demanding and require careful parameter tuning and scenario-specific constraints to yield reliable results.
To overcome limitations of physics-based methods and combine their advantages with the power of learning-based approaches, hybrid models [224, 225, 226, 227] have been proposed by the community. Kim et al. [224] proposed a hybrid method that uses the Reciprocal Velocity Obstacles (RVO) framework to predict pedestrian trajectories in dynamic environments, enhanced with Ensemble Kalman Filtering (EnKF) and Expectation-Maximization (EM) for real-time learning and parameter refinement. Ma et al. [225] introduced a deep latent variable model built on a Wasserstein auto-encoder, enabling the automatic satisfaction of kinematic constraints. Yue et al. [226] presented Neural Social Physics (NSP), a Neural Differential Equation model that integrates an explicit physics-based framework as an inductive bias, paired with deep neural networks for system parameter estimation and stochastic dynamics modelling.
@@ 3.3 Machine Learning-based Methods
Machine Learning-based methods learn motion dynamics from driving data by fitting function approximators. Figure 9 provides an overview of two key ML techniques discussed in this section: Gaussian Processes (GPs) and Hidden Markov Models (HMMs). Additionally, the section discusses prediction models based on Gaussian Mixture Models (GMMs), which combine multiple GPs, and Dynamic Bayesian Networks (DBNs), which capture interactions between agents. Each method is discussed in dedicated subsections, elaborating on the theoretical foundations, their application to prediction tasks, and relevant studies available in the literature.
@@ 3.3.2 Hidden Markov Models
Rathore et al. [238] proposed a hybrid framework, Traj-clusiVAT, combining trajectory clustering and Markov chain modelling. The clustering algorithm partitions trajectories into representative clusters, and Markov models are trained for prediction. Li et al. [239] proposed an enhanced HMM for behavior prediction, with key improvements including random initial parameter generation to ensure reliability, parameter learning through the Baum-Welch algorithm, and prediction via the Viterbi algorithm combined with Bayesian decision theory. On Continuous HMMs, Ren et al. [240] proposed a framework for behavior recognition. The CHMM is employed for early detection of lane-changing behavior by analyzing vehicle motion patterns and contextual traffic features.
@@ 3.3.4 Dynamic Bayesian Networks
3.3.5 Strengths and Limitations of ML Methods
The discussed ML-based methods offer transparent structure enabling straightforward interpretation and debugging, as the models rely on well-defined mathematical components such as transition probabilities and covariance functions. The methods also provide uncertainty estimates and can perform well with limited training data. However, their limited parameterization restricts their ability to model complex, high-dimensional behaviors common in real-world traffic. DBNs, which can model interactions, struggle to scale with the number of agents or observations and require manual design of model structure and features. Unlike deep learning models, they cannot learn hierarchical representations, reducing their adaptability to new environments or interaction patterns.
@@ 3.4 Deep Learning-based Methods
Deep Learning-based methods excel in learning motion dynamics from driving data spanning different architectures. Figure 10 illustrates the main DL models reviewed in this section. Each approach is detailed in a separate subsection, with methods summarized in tables that highlight the key features of the architectures.
3.4.1 RNNs and LSTMs
@@
Sequential models such as RNNs and LSTMs assume that motion evolves through temporally causal dependencies. By maintaining hidden states across time steps, they effectively model dynamic behavior and long-term temporal patterns for trajectory prediction. Focusing on model design, earlier studies primarily utilized a single recurrent model to map past trajectories to future paths [285, 97, 102]. However, more recent works employ multiple groups of recurrent models to address different aspects of motion [286, 253, 22], as well as stacked LSTMs and convolutional LSTMs. Dai et al. [22] proposed a model, where one group of LSTMs models individual vehicle motions, while another group captures interactions between agents. The architectural approach for stacked LSTMs involves hierarchically organizing recurrent networks into layers, as seen in [287, 248]. Tables 3 and 4 provide a summary of architectures in the literature that rely on recurrent models.
@@ 3.4.2 CNNs
Convolution Neural Networks for trajectory prediction are categorized into two groups in the literature: those using image-based input and Temporal Convolution Networks (TCNs). Image-based inputs can come from various perspectives, including BEV, vehicle-mounted cameras, or other sensor modalities. BEV images or aerial photographs offer a top-down perspective that effectively captures spatial relationships and road structures. CNN-based models [69, 61, 295] primarily focus on the feature extraction stage to capture road geometries, traffic agent positions, and environmental context. The second category involves Temporal Convolution Networks (TCNs), which apply convolution layers directly to trajectory sequences to model temporal dependencies. Unlike RNNs, TCNs operate in parallel over the entire sequence, incorporating causal convolutions to preserve temporal order and dilated convolutions to extend the receptive field, capturing long-term dependencies. TCNs are designed to maintain stable gradients, avoid information leakage, and achieve computational efficiency for long sequences. One example is the framework proposed by Nikhil and Morris [86], which preserves spatio-temporal consistency through multiple convolution layers while leveraging parallelism for efficiency. Table 5 provides a summary of CNN-based architectures for trajectory prediction.
@@ 3.4.3 Autoencoders
In trajectory prediction, autoencoders encode past trajectories into compact latent representations, which are then decoded into future positions. These models excel at capturing complex temporal dependencies and can function as standalone predictors or feature extractors integrated into larger architectures. Variational Autoencoders (VAEs) extend this approach by modeling uncertainty, generating a latent distribution instead of fixed representations. By sampling from this distribution, VAEs generate diverse and multimodal future trajectories, modelling the inherent uncertainty in trajectory forecasting tasks. Table 6 provides a summary of autoencoder architectures available in the literature, grouped by deterministic or stochastic type.
@@ 3.4.4 GANs
Generative Adversarial Networks (GANs) also found their application in trajectory prediction, leveraging their ability to model complex, multimodal data distributions. In this context, the generator ($G$) predicts plausible future trajectories conditioned on past trajectories, while the discriminator ($D$) evaluates these predictions against real trajectory data to guide the generator’s learning. This adversarial training setup enables GANs to generate diverse and realistic trajectories that account for the inherent uncertainty in future motion. A key advantage of GAN-based models is their capacity to address the multimodal nature of trajectory prediction, where multiple plausible futures exist for a given past trajectory. This is achieved by conditioning $G$ on contextual information such as social interactions, environmental constraints, or road semantics, allowing it to adapt predictions to dynamic and complex scenarios. Table 7 provides a summary of GAN architectures in the literature, detailing the design of $G$ and $D$.
@@ 3.4.6 Attention and Transformers
Attention mechanisms enhance prediction models by allowing them to focus on the most relevant spatial and temporal features. Spatial attention mechanisms assign importance weights to neighboring agents computed at each timestamp, based on features such as relative distance, velocity, and heading, which influence the target agent’s trajectory. Temporal attention processes sequences of historical states, assigning weights to time steps that have the most significant impact on future predictions. This is often integrated with recurrent architectures, where the hidden states from prior time steps are passed through an attention module to refine the context vector for trajectory prediction. Advanced implementations combine spatial and temporal attention in multi-head architectures, enabling the model to process multiple attention patterns in parallel, effectively handling heterogeneous agents in dense environments. Table 11 provides a summary of reviewed models relying on attention mechanisms.
@@
3.4.7 Strengths and Limitations of DL Methods
Deep learning-based trajectory prediction models exhibit varying strengths and limitations depending on their architectural design. LSTM-based models effectively capture temporal dependencies and can be trained in an autoregressive mode, similar to inference mode. They converge reliably but lack spatial reasoning and struggle with long-range temporal dependencies. CNN-based models efficiently extract spatial features from image-based inputs such as BEV or rasterized maps and enable fast inference. In Autoencoders, their variational forms support multimodal prediction by modeling uncertainty, but their performance depends heavily on the latent space structure and they may produce unrealistic trajectories if not properly trained. GAN-based models can learn complex trajectory distributions and generate diverse outputs, yet they require large training datasets, can be challenging to stabilize, and may not generalize well under distribution shifts when run in out-of-distribution scenarios. GNN-based models excel in dense interaction scenarios by explicitly modeling agent-agent relationships, but they add unnecessary complexity in sparse environments such as highway driving, where interactions are minimal. In this case, rather than model limitations, it becomes a matter of appropriate design choice. Dynamic graph-based GNNs depend on consistent object IDs for temporal tracking and can suffer from tracking errors. Transformers, due to their ability to capture long-term dependencies, are expected to excel in long-horizon prediction. However, they are highly sensitive to training stability, require velocity-based inputs (e.g., $v_{x},v_{y}$) for convergence (which is beneficial for all DL models), and demand extensive tuning and warm-up for optimal performance.
@@ 3.5 Reinforcement Learning-based Methods
Reinforcement Learning-based methods generate predictions by interpreting agent behaviour as a high-dimensional policy optimization problem. RL methods learn through a reward system to generate trajectories that maximize long-term objectives. This section reviews two main categories of methods: Inverse Reinforcement Learning and Imitation Learning.
3.5.1 Inverse Reinforcement Learning